OpenStudy (anonymous):

square root of 3 open parentheses square root of 6 plus square root of 3 close parentheses plus open parentheses square root of 8 end root minus 5

OpenStudy (mertsj):

\[\sqrt{3}(\sqrt6+\sqrt{3})+(\sqrt{8}-5)=\sqrt{18}+3+\sqrt{8}-\sqrt{5}=3\sqrt{2}+3+2\sqrt{2}-\sqrt{5}=\]

OpenStudy (mertsj):

\[5\sqrt{2}+3-\sqrt{5}\]